Electric field in a regionregion is E=7x^2 -2x +2. Find the electric potential at sime point in this region.

Grade:12

3 Answers

Eshan
5 years ago
The electric potential as a function ofxwill be given as
V=-\int E.dx=\int(7x^2-2x+2)dx
=\dfrac{7x^3}{3}-x^2+2x+constant

To exactly know the value of the constant, we need to know the potential at some reference point.
Gitanjali Rout
5 years ago
To exactly know the value of the constant, we need to know the potential at some reference point.=\dfrac{7x^3}{3}-x^2+2x+constantV=-\int E.dx=\int(7x^2-2x+2)dxwill be given asxe electric potential as a function of
Forum Team
5 years ago
 
 
Gitanjali, please stop copy pasting answers posted by other users. You will not be rewarded for these points and your account will be blocked due to this activity as it is against forum guidelines. Please check the guidelines here – Forum point Policy – https://www.askiitians.com/forum-point-policy/
 
Thanks
Forum Team